Fermi National Laboratory
 

Accelerator Update for May 29, 2009 - June 1, 2009

 

The Integrated Luminosity 5/25/09 to 6/1/09 was 51.59 pb-1.

Friday May 29

The day shift began with Tevatron (TeV) studies following the resolution of a TeV beam permit problem, with Operations monitoring the Recycler stash (278.8E10), with the Antiproton Source stacking (24.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, NuMI, and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

The H- Source experienced high voltage regulation problems and the water resistor needed cleaning before it returned to normal operation.

Operations put the TeV into HEP shot setup and established store 7095 at 5:45 PM with an average initial luminosity of 303E30.

Linac quadrupole power supplies 110-113 tripped off requiring attention from Linac personnel. The mains breaker was reset and beam was reestablished.

Pbar transferred antiprotons to the Recycler when appropriate.

Saturday May 30

The midnight shift began with Operations monitoring store 7095 (132.64E30) and the stash (190.6E10), with Pbar stacking (23.6E10), and with NuMI and MiniBooNE taking beam.

Operations contacted Linac personnel in hopes of recovering the H- Source. Efforts were successful and the intensity was back up by 3 AM.

Operations established beam to MTest.

Operations terminated store 7095 at 8:30 AM.

Operations put the TeV into HEP shot setup and established store 7096 at 10 AM with an average initial luminosity of 294E30.

Operations disabled the MTest beam at 6 PM.

Operations terminated store 7096 at 10:15 PM and put the TeV into HEP shot setup.

Pbar transferred antiprotons to the Recycler when appropriate.

Sunday May 31

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established store 7097 at 12:10 AM with an average initial luminosity of 278E30.

The MI-10 injection kicker #1 required a local reset.

Operations established beam to MTest.

At 9:49 AM, store 7097 quenched at TeV sector B0. A B0 quadrupole power supply (B0QT3) took a 4-amp walk and caused the quench. Operations called in EE Support personnel to replace the unit with a spare.

Pbar VSA problems prompted Operations to call in Pbar personnel to investigate. The problem eventually cleared and normal operations resumed.

A false Booster beam valve indication (the page said the valves were moving) caused some downtime while the issue was resolved.

The H- Source output decreased throughout most of the day shift. Operations called in Linac personnel to assist in stabilizing the output.

Operators replaced a MI correction magnet amplifier (V407).

Operations reported problems with LBSEQ during TeV turn-on that resulted in a D0 low beta quench. A bad 10-volt VFC card was replaced following the D0 quench.

Delays in recovering the TeV sector C4 cryo system continued through the rest of the evening shift.

Monday June 1

The midnight shift began with the TeV cryo system recovering at sector C4, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations began the proton intensity reduction program in preparation for the 11-week shutdown beginning on June 15 th.

A valve on the C4 wet engine was found to be leaking and prompted a change out by CHL technicians. The newly installed valve was found to be leaky as well after having continued problems in recovering the house. The valve was changed out a second time and the C4 cryo system was ready for HEP operations by 3 AM.

Operations began the TeV check out and turn on and a ramp was established by 5:45 AM.

Operations put the TeV through a dry squeeze and then checked orbits.

Operations put the TeV into HEP shot set-up at 7 AM.

Pbar transferred antiprotons to the Recycler when appropriate.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/01/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory